OpenStudy (anonymous):

Which statement describes the translation of y= –5(x2 + 2) – 3 from standard position?

OpenStudy (amistre64):

depends on what the statements are

Directrix (directrix):

@BMI229 --> What are the options (statements) from which to choose?

OpenStudy (anonymous):

y 4 y 4 y 3 y 3

OpenStudy (anonymous):

y < 4 y > 4 y >3 y <3 and are all have the line underneath

OpenStudy (amistre64):

also chk to make sure youve posted the correct equation y = -5(x+2)^2 - 3 looks better to me, but i cant see your material

OpenStudy (anonymous):

Which statement describes the translation of y= –5(x2 + 2) – 3 from standard position? Moved up and to the right. Moved up and to the left. Moved down and to the left. Moved down and to the right.

OpenStudy (anonymous):

thats the problem sorry took answers off another questions

OpenStudy (amistre64):

id have to assume its asking for the movement of say the vertex from (0,0) to (h,k)

OpenStudy (amistre64):

given a vertex form as: y = a(x-h)^2 + k

OpenStudy (anonymous):

???

OpenStudy (anonymous):

So which it? Moved up and to the right. Moved up and to the left. Moved down and to the left. Moved down and to the right.

OpenStudy (amistre64):

this site is not a free answering service, it is a site to help you study the material. If you have questions about what I have presented, then ask about what is confusing you please.

OpenStudy (anonymous):

I don't get it?

OpenStudy (anonymous):

what are the directions of the quadratic its for a study guide

OpenStudy (amistre64):

i have to make a few assumptions here since I do not have the problem you are working on and I believe that you have made some typos. a standard position of a graph is when its vertex is at the origin (0,0) there are hints given in the equation that relate to how the graph has been moved from the origin

OpenStudy (amistre64):

given a standard position quadratic equation: y = x^2 if it has been moved away from the origin it will exhibit the property of: y = (x-h)^2+k the values of h and k describe the movements.

OpenStudy (anonymous):

h is what and k is what?

OpenStudy (amistre64):

you are given: y = -5(x+2)^2 - 3 or rewriten as: y = -5(x-(-2))^2 + (-3) h k how would we move from the point (0,0) to the point (-2,-3) ?